OpenStudy (anonymous):

MEDAL! A system of equations is shown below: x + y = 3 2x - y = 6 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

So you're saying that the solution to the system of equations is (9, 6), @15238?

OpenStudy (anonymous):

Because 9 + 6 does not equal 3. :/

OpenStudy (anonymous):

But 2 * 9 - -6 = 24.

OpenStudy (anonymous):

In the second equation in the system of equations, we are subtracting, not adding.

OpenStudy (anonymous):

Oh, I made a mistake in my head. x = 3 and y= 0

OpenStudy (anonymous):

2x-6 = 3 - x 3x = 9 x = 3

OpenStudy (anonymous):

Yep, that checks out! I can't believe I didn't think of that! I tried like every other possible solution. Lol! Thank you so much, @15238! :)

OpenStudy (anonymous):

Yeah np, just isolate y in each equation then set them equal to each other